OpenStudy (anonymous):

differentiate \tan ^{-1}\left( \frac{ \sin2x - \cos2x }{ \sin2x +\cos2x }\right) with relation to x

OpenStudy (anonymous):

\[\tan ^{-1}\left( \frac{ \sin2x - \cos2x }{ \sin2x +\cos2x }\right)\]

OpenStudy (anonymous):

y = (the whole arctan thingy)

OpenStudy (shamil98):

would it not be whatever (arctan)' * the fraction thing ? if let the fraction thing = u \[\tan^-1(u) = \frac{ 1 }{ 1+u^2 }\]

OpenStudy (shamil98):

\[(\tan^{-1} u)' \] my bad

OpenStudy (shamil98):

I'm thinking it's either that or it involves the chain rule?..

OpenStudy (anonymous):

if you use the chain rule, you can get the answer but the steps are really ugly.... i dont think the examiner would accept the easy way.... there was some method which used the tan 2x identity

zepdrix (zepdrix):

Hmmmm I can't think of what identity that would be..

OpenStudy (shamil98):

me niether..

OpenStudy (anonymous):

i have solved it using the example zepdrix showed me. it works but isnt there a clean and easy way..? i remember there was

OpenStudy (anonymous):

anyway, thnx guys

zepdrix (zepdrix):

Oh oh oh I think I got it.

zepdrix (zepdrix):

Hold on lemme check my steps really quick.

OpenStudy (anonymous):

geeez i can´t wait!!

zepdrix (zepdrix):

Omg............ stupid browser just erased all of my text :(

zepdrix (zepdrix):

omggggggggggggggggggggggggggggggggggggggggggg

OpenStudy (anonymous):

noooooooooooooooooooooooOOOOOOO!

zepdrix (zepdrix):

I made a mistake somewhere anyway... I thought it simplified to tan4x, but I made a lil error, I think it simplifies to sec4x-tan4x. I'm not sure if that will work for us.. hmm

OpenStudy (anonymous):

did you divide the num and denom by cos2x first?

zepdrix (zepdrix):

I multiplied the numerator and denominator by the conjugate of the denominator. Then expanded out the square in the numerator, and applied the Sine and Cosine Double Angle Identities. Grrr I don't wanna write it out again +_+ Divide by cos2x? is that a good approach maybe? Hmm

OpenStudy (anonymous):

i tried using the conjugate too but i wasn´t getting anywhere.. i´ll try once more

zepdrix (zepdrix):

\[\Large\bf\sf =\frac{\tan2x-1}{\tan2x+1}\]So dividing by cos2x gives us this or something? hmm

OpenStudy (anonymous):

yepp

OpenStudy (anonymous):

ok, i got tan4x - sec4x

zepdrix (zepdrix):

I have an idea, I'm gonna try using the tangent sum/difference identity. \[\Large\bf\sf \tan\left(\alpha-\beta\right)\quad=\quad \frac{\tan \alpha-\tan \beta}{1+\tan \alpha \tan \beta}\]

OpenStudy (anonymous):

dude this is not tan this is ARCtan, so r u sure you can use that?

zepdrix (zepdrix):

\[\large\bf\sf \tan\left(2x-\frac{\pi}{4}\right)\quad=\quad \frac{\tan2x-\tan\frac{\pi}{4}}{1+\tan2x \tan\frac{\pi}{4}}\quad=\quad \frac{\tan2x-1}{\tan2x+1}\]Ooo I think that works, doesn't it?

zepdrix (zepdrix):

This is just to deal with the inside part still.

OpenStudy (anonymous):

yep!! i think that´s it woooT!

OpenStudy (anonymous):

cool!! thanks a lot zep!

zepdrix (zepdrix):

\[\Large\bf\sf \arctan\left(\tan(2x-\pi/4\right)\quad=\quad 2x-\pi/4\]Something like that? :D yayyy team lol

OpenStudy (anonymous):

i see, the arctan and tan cancel out and then you differentiate 2x - pi/4

OpenStudy (anonymous):

now this is what i call differentiating with swag!

zepdrix (zepdrix):

Ooo nice we got the correct answer. I checked it in Wolfram just now using the long messy one. Derivative = 2